OpenStudy (anonymous):

find limit as x approaches 0: tan(x)/ (7x + sin(x))

myininaya (myininaya):

Do you know l'hospital's rule? Or do you recall that \[\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=1\]

OpenStudy (anonymous):

i recall the second one. i'd like to know how to do it with both

myininaya (myininaya):

Well you have x->0 if you plug in 0 and you get an indeterminate form such as 0/0 or inf/inf then you can do l'hospital's rule derivative of top/derivative of bottom

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\tan(x)}{7x+\sin(x)}=\lim_{x \rightarrow 0}=\frac{(\tan(x))'}{(7x+\sin(x))'}\]

myininaya (myininaya):

OR! \[\lim_{x \rightarrow 0}\frac{\frac{x \sin(x)}{x \cos(x)}}{7x+x \frac{\sin(x)}{x}}\]

myininaya (myininaya):

I wrote tan(x) as sin(x)/cos(x) Putting it in terms of sin(x) and cos(x) helps since we know that one limit property plus another which is \[\lim_{x \rightarrow 0}\frac{\cos(x)-1}{x}=0\] which probably won't be used here so we have \[\lim_{x \rightarrow 0}\frac{\frac{x}{\cos(x)} \frac{\sin(x)}{x}}{7x+x \frac{\sin(x)}{x}} \] Where does sin(x)/x go again as x goes to 0?

myininaya (myininaya):

Do you have any questions?

myininaya (myininaya):

Can you answer my question?

OpenStudy (anonymous):

the brackets are confusing. i changed tanx to sinx/cosx, which then gives me sinx/ (cosx(7x+sinx). but i'm not sure where to go from there since it's still 0/0

myininaya (myininaya):

Did you see I multiply sin(x)/cos(x) by x/x and sin(x) by x/x

OpenStudy (anonymous):

huh? i couldn't understand the notation. and why?

myininaya (myininaya):

the notation? the fractions?

OpenStudy (anonymous):

yeah. it's in bracket form. i dunno. i got to 2x/ (14cosx + sin2x) and i'm stuck

myininaya (myininaya):

is the code not showing up as fractions?